OpenStudy (jaylelile):

WILL MEDAL What is the multiplicative rate of change for the exponential function f(x) = 2()–x?

OpenStudy (jaylelile):

please help! Ginny is studying a population of frogs. She determines that the population is decreasing at an average rate of 3% per year. When she began her study, the frog population was estimated at 1,200. Which function represents the frog population after x years? f(x) = 1,200(1.03)x f(x) = 1,200(0.03)x f(x) = 1,200(0.97)x f(x) = 1,200(0.97)x

OpenStudy (jaylelile):

@jpaulkai ?

OpenStudy (anonymous):

Well if 1.00 represents 100%, what would equal 3%? (HINT: Move the decimal around)

OpenStudy (anonymous):

Also, why are the last 2 answers the same?

OpenStudy (jaylelile):

I have no idea lol

OpenStudy (anonymous):

Okay well lets use the process of elimination. If 1.00 is 100%, then 1.03 is 103% - which is incorrect. That also means that 0.97 is 97%, which is also incorrect. Leaving us with one other option since the last two options are identical.

OpenStudy (anonymous):

THE ANSWER IS THE SECOND ONE SO... (B)